# Data-Effective Learning: A Comprehensive Medical Benchmark

### v1

@@
Abstract
Data-effective learning aims to use data in the most impactful way to train AI models, which involves strategies that focus on data quality rather than quantity, ensuring the data used for training has high informational value. Data-effective learning plays a profound role in accelerating AI training, reducing computational costs, and saving data storage, which is very important as the volume of medical data in recent years has grown beyond many people’s expectations. However, due to the lack of standards and comprehensive benchmark, research on medical data-effective learning is poorly studied. To address this gap, our paper introduces a comprehensive benchmark specifically for evaluating data-effective learning in the medical field. This benchmark includes a dataset with millions of data samples from 31 medical centers (DataDEL), a baseline method for comparison (MedDEL), and a new evaluation metric (NormDEL) to objectively measure data-effective learning performance. Our extensive experimental results show the baseline MedDEL can achieve performance comparable to the original large dataset with only 5% of the data. Establishing such an open data-effective learning benchmark is crucial for the medical AI research community because it facilitates efficient data use, promotes collaborative breakthroughs, and fosters the development of cost-effective, scalable, and impactful healthcare solutions. The project can be accessed at https://github.com/shadow2469/Data-Effective-Learning-A-Comprehensive-Medical-Benchmark.git.
Figure 1: (a) Data-effective learning aims to obtain a compact small dataset from a large-scale dataset, but the two datasets have similar effects on AI model training. (b) Demonstration of our comprehensive benchmark for data-effective learning. The benchmark includes a dataset of millions of data samples from 31 medical centers (DataDEL), a baseline method for comparison (MedDEL), and a new evaluation metric (NormDEL).
@@ Abstract
Model Pre-training Gastrovision Jha et al. (2023) Detection and classification 2 None 8,000 GI
Hyper-Kvasir Borgli et al. (2020) Classification 1 374 110,079 GI
Kvasir-Capsule Smedsrud et al. (2021) Classification 1 117 47,238 GI
@@
Model downstream testing CVC-12k(CVC-ClinicDB) Bernal et al. (2015) Segmentation 1 None 612 polyp
CVC-300 Vázquez et al. (2017) Segmentation Not mentioned None 60 polyp
CVC-ColonDB Bernal et al. (2012) Segmentation Not mentioned None 380 polyp
EAD2019 Ali et al. (2019) Endoscopic artifacts 6 None 2,991 cancers
EDD2020 Ali et al. (2020) Segmentation, detection, and localization 4 None 380 polyp
ETIS Silva et al. (2014) Segmentation Not mentioned None 196 polyp
ImageCLEFmed Hersh et al. (2009) Polyp segmentation 1 None 66,662 polyp
Kvasir-Instrument Jha et al. (2021a) Segmentation, detection, and localization 1 None 590 polyp
Kvasir-SEG Jha et al. (2020) Segmentation, detection, and localization 4 None 1,000 polyp
Kvasir-Sessile Jha et al. (2021b) Segmentation, detection, and localization 4 None 196 polyp
PolypGen2021 Ali et al. (2021) Polyp segmentation 6 None 3,762 polyp
Table 1: Overview of the proposed DataDEL in our benchmark. We integrate 31 medical centers, including data on the scale of millions, to build a high-quality, large-scale, multi-disease comprehensive dataset, aiming to provide researchers with a better data platform.
1 Introduction
@@
Here, we introduce the concept of Data-Effective Learning, which plays a significant role in the field of medical data, accelerating model training and reducing storage burden in the era of big data (Figure 1(a)). The global endoscopy surgery market is undergoing significant development, with an estimated daily addition of 22,546,800,000 video frames  end (2023). However, a substantial presence of disruptive and invalid data significantly hampers training efficiency and occupies a considerable amount of storage space. Therefore, achieving data-effective in endoscopy datasets holds the following special advantages:
@@
  • Storage Savings: Assuming the use of traditional high-definition endoscopes (1080p) He et al. (2021b), the daily uncompressed endoscopy examination videos would require 12,756,493 TB (about 13.06 exabytes) of storage space. Condensing endoscopy datasets can result in substantial storage savings.
@@
  • Enhanced Model Efficiency: Borgli et al. (2020) released the largest digestive system image dataset (Hyper-Kvasir) to date and showed that upon analyzing data sources, over 90% of video frames consist of disruptive and invalid data. Core critical data comprises only 2% of the entire dataset  Jha et al. (2020, 2021b). Efficient utilization of core critical data can significantly improve the efficiency of model training, in order to achieve rapid convergence.
@@
  • Computational Resource Savings: With the use of a single RTX 3090 graphics card and the VGG16 model Kim et al. (2016) (approximately 138 million parameters) for 32-bit floating-point (FP32) calculations, processing 325.6 images per second is achievable Kwasny et al. (2020). In this configuration, training on the daily added video frames would require 19,200 hours. Utilizing only core critical data could save nearly 18,816 hours without compromising precision.
Figure 2: Pipeline of the baseline method (MedDEL) for data-effective learning in our benchmark. It illustrates effective removal of disruptive and invalid data from the dataset, aiming to save storage space and computational resources while enhancing model efficiency.
@@
The issue of data-effective learning is not exclusive to endoscopy datasets; it is also evident in other types of medical datasets. With the rapid expansion of future medical data, efficiently handling medical datasets is the next crucial research problem in data-driven learning methods.
@@
In recent years, there has been relevant research on data-effective learning in natural image datasets such as adversarial samples Feinman et al. (2017), and data biases Wang et al. (2019), aiming to enhance the robustness and generalization capabilities of deep learning models. These works in natural datasets have shown preliminary effectiveness, indicating significant potential advantages such as data storage, computational resource savings, and efficient model training. However, despite the exponential growth of medical datasets, there is currently a lack of such exploration in the medical field Xiao et al. (2023). This is attributed to the absence of relevant benchmarks, encompassing unified datasets, baseline methods, and comprehensive evaluation metrics, providing the academic community with a basis for developing and comparing advanced methods.
@@
In addition to the absence of benchmark, achieving data-effective in the field of endoscopy remains a huge challenge Sharma et al. (2021) and is worth exploring in depth. Traditional data-effective techniques may struggle to work effectively with datasets that have high-dimensional features or sparse data. In endoscopy datasets, the challenge lies more in identifying subtle differences in images, which can be crucial for medical diagnosis. Therefore, accurately determining the similarity of images in endoscopy datasets is a top priority in research. Maintaining data quality and integrity is a key issue in data-effective learning, as it directly affects the usability and accuracy of the dataset.
@@
2 Related Work
In the natural image domain, benchmarks for assessing data-effective learning typically involve various techniques such as structural similarity index, convolutional neural networks, and local feature descriptors. These methods are evaluated on widely used image datasets like ImageNet and CIFAR, or specialized datasets. PXDedup Xie et al. (2021) explores the issue of data-effective learning in JPEG images, pointing out that traditional binary stream-based techniques do not work well for compressed JPEG images.  Chen et al. (2013) introduces a high-precision image data effective method that identifies and eliminates duplicate images through feature extraction and high-dimensional indexing.  Chen et al. (2012) uses wavelet decomposition to extract feature vectors from images and calculates the Manhattan distance to determine image similarity, thus achieving the detection and removal of duplicate images. However, when handling datasets with high-dimensional features or sparse data, traditional data-effective techniques might not work effectively. MFDedup Zou et al. (2022) surpasses existing technologies in improving data reduction rates and recovery throughput, while also reducing the cost of garbage collection. Storage and memory capacity can also become limiting factors.  Harnik et al. (2012) discusses how to accurately estimate the data reduction ratio achieved through data effective and compression techniques for specific datasets. Hash methods indeed have a wide application in data effective processing. The CE-Dedup Li et al. (2021) framework combines hash-based image data-effective techniques with deep learning image classification tasks. By adjusting the deduplication threshold, effectively balances the trade-off between data reduction rate and model accuracy.  Liu et al. (2016) proposes a deep supervised hash algorithm for image retrieval by providing the model with pairs of images labeled as similar/dissimilar.
@@ 2 Related Work
3 Data-Effective Learning Medical Benchmark
In this section, we introduce the definition of benchmark tasks. Data-effective learning refers to the practice of using a limited amount of data for the pre-training phase. Therefore, this benchmark encourages researchers to develop advanced data-effective learning methods to generate a compact version of the originally collected large-scale dataset, thereby obtaining a compact small-scale new dataset for pre-training foundational models.
In the era of big data, large models often require massive pre-training data. However, our perspective challenges this conventional thinking, and we have undertaken the following work to explore this issue in depth.
@@ 3 Data-Effective Learning Medical Benchmark
3.1 The benchmark dataset: DataDEL
Given the urgent demand for a comprehensive benchmark in the medical field, we are facing the pressing task of integrating diverse large datasets. Currently, prevailing challenges within existing medical datasets encompass issues like the uniformity of data sources, the standardization of task execution, the limited scope of covered disease types, imbalances in dataset categories, and the uniformity of modalities.
@@ 3.1 The benchmark dataset: DataDEL
3.2 The benchmark baseline: MedDEL
We introduce a benchmark baseline method (MedDEL) in the endoscopic medical field, which is based on the principles of the SemDeDup method Abbas et al. (2023). Endoscopic data holds significant importance in medical diagnosis and treatment. However, due to various factors such as organ morphological changes Smith and Nanci (2003), lighting conditions Clancy et al. (2012), and noise interference Gulenko et al. (2022), the analysis of this data becomes complex and challenging.
@@ 3.2 The benchmark baseline: MedDEL
3:     for $cluster_{i}$in $clusters$ do
4:        for $j=1$ to $Size(cluster_{i})$ do
5:           if $dis(F_{x,j},centers_{i})>\epsilon$ then
@@
8:              for $k=j$ to $Size(cluster_{i})$ do
9:                 if $Similarity(F_{j},F_{k})>\eta$ then
10:                    Del $min(dis(F_{j},center_{i}),dis(F_{k},center_{i}))$
11:                 end if
12:              end for
13:           end if
14:        end for
15:     end for
16:     Increment iteration counter: $t\leftarrow t+1$
17:  end while
Specifically, the MedDEL method takes an image $I$ as an $i^{th}$ example, the encoder of Vision Transformer (ViT) Dosovitskiy et al. (2020) extracts deep features from the image, ultimately producing a 768-dimensional feature output. This output serves as our information embedding. The high-dimensional representation not only captures the semantic information of the image more effectively but also addresses semantic repetitiveness issues that are challenging to resolve in low-dimensional spaces. Considering the encoding process of image $I$ through the ViT model, we can represent it with the following formula:
$F_{x}=ViT(I)$ (1)
Where $x$ represents the final layer of the $ViT$ encoder.
However, by introducing the K-means algorithm Macqueen (1967), the overall computational complexity is reduced from $O(n^{2})$ to $O(N^{2}/k)$, where $k$ is the number of clusters. After applying the K-means algorithm, we obtain $k$ clusters. Assuming that the feature $F(x,i)$ belongs to the $k$ cluster, we can derive the following $k$ sequences:
$C_{k}$ =$=\{F(x,i)\mid i\in N^{*},F(x,i)\in cluster_{k}\}$ (2)
@@
Table 2: Experimental setting for the parameter of $\eta$ in Algorithm  1 that controls the data remaining ratio. We set several different ratios to fully assess the performance of MedDEL under various proportions of remaining data. Additionally, we also calculate the number of training epochs required for different ratios of data at the same computational power for fair comparison.
@@
Figure 3: Demonstration of the feasibility of data-effective learning. We compared the performance differences between using only 5% of the pre-training data (in red) and using 100% of the data (in black) in 8 datasets. The results indicate that using only 5% of the pretraining data can achieve results comparable to using 100% of the pretraining data, which fully demonstrates the validity of the MedDEL method.
@@
Table 3: Demonstration of the rationality of NormDEL. The table shows the performance of different proportions of pre-trained data on mIoU ($\uparrow$) as well as NormDEL ($\uparrow$) in eight different datasets. The results indicate that the performance of mIoU is almost comparable in different scales, suggesting that it primarily measures performance without considering the scale of the pretraining data used. In contrast, NormDEL incorporates the factor of data scale, and can evaluate data compactness even with only 5% of the data used. This demonstrates the rationality of the NormDEL method, as it not only assesses performance but also effectively utilizes the data scale.
3.3 The benchmark evaluation metric: NormDEL
Currently, there is no comprehensive objective metric for evaluating the performance of data-effective learning algorithms, where comprehensive refers to integrating both the test accuracy of downstream tasks and the compactness of pre-training data. Hence, We propose a new data-effective metric, the Normalized Delete Index (NormDEL). Our objective is that within the same task, if a model can achieve the same level of performance in downstream tasks using fewer pre-training data, its data-effective performance should be considered superior. Taking the commonly used segmentation task in endoscopy as an example, where the general metric for measuring performance is often mIoU (mean Intersection over Union), we can formulate the following equation:
$DEL=mIoU\cdot e^{-\alpha\cdot R}$ (4)

### v3

# A Medical Data-Effective Learning Benchmark for Highly Efficient Pre-training of Foundation Models

@@
Abstract.
Foundation models, pre-trained on massive datasets, have achieved unprecedented generalizability. However, is it truly necessary to involve such vast amounts of data in pre-training, consuming extensive computational resources? This paper introduces data-effective learning, aiming to use data in the most impactful way to pre-train foundation models. This involves strategies that focus on data quality rather than quantity, ensuring the data used for training has high informational value. Data-effective learning plays a profound role in accelerating foundation model training, reducing computational costs, and saving data storage, which is very important as the volume of medical data in recent years has grown beyond many people’s expectations. However, due to the lack of standards and comprehensive benchmark, research on medical data-effective learning is poorly studied. To address this gap, our paper introduces a comprehensive benchmark specifically for evaluating data-effective learning in the medical field. This benchmark includes a dataset with millions of data samples from 31 medical centers (DataDEL), a baseline method for comparison (MedDEL), and a new evaluation metric (NormDEL) to objectively measure data-effective learning performance. Our extensive experimental results show the baseline MedDEL can achieve performance comparable to the original large dataset with only 5% of the data. Establishing such an open data-effective learning benchmark is crucial for the medical foundation model research community because it facilitates efficient data use, promotes collaborative breakthroughs, and fosters the development of cost-effective, scalable, and impactful healthcare solutions. The benchmark can be accessed at GitHub Repository.
Medical benchmark, Data-Effective learning, Endoscopic Image Processing, Foundation Model
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681313isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision
Table 1. Overview of the proposed DataDEL in our benchmark. We integrate 31 medical centers, including data on the scale of millions, to build a high-quality, large-scale, multi-disease comprehensive dataset, aiming to provide researchers with a better data platform.
@@ Abstract.
Model Pre-training Gastrovision (Jha et al., 2023) Detection and classification 2 None 8,000 GI
Hyper-Kvasir (Borgli et al., 2020) Classification 1 374 110,079 GI
Kvasir-Capsule (Smedsrud et al., 2021) Classification 1 117 47,238 GI
@@
Model downstream testing CVC-12k (CVC-ClinicDB) (Bernal et al., 2015) Segmentation 1 None 612 polyp
CVC-300 (Vázquez et al., 2017) Segmentation Not mentioned None 60 polyp
CVC-ColonDB (Bernal et al., 2012) Segmentation Not mentioned None 380 polyp
EAD2019 (Ali et al., 2019) Endoscopic artifacts 6 None 2,991 cancers
EDD2020 (Ali et al., 2020) Segmentation, detection, and localization 4 None 380 polyp
ETIS (Silva et al., 2014) Segmentation Not mentioned None 196 polyp
ImageCLEFmed (Hersh et al., 2009) Polyp segmentation 1 None 66,662 polyp
Kvasir-Instrument (Jha et al., 2021a) Segmentation, detection, and localization 1 None 590 polyp
Kvasir-SEG (Jha et al., 2020) Segmentation, detection, and localization 4 None 1,000 polyp
Kvasir-Sessile (Jha et al., 2021b) Segmentation, detection, and localization 4 None 196 polyp
PolypGen2021 (Ali et al., 2021) Polyp segmentation 6 None 3,762 polyp
Figure 1. Data-Effective Learning (DEL) enables more efficient pre-training of foundational models. (a) Data-effective learning aims to obtain a compact small dataset from a large-scale pre-training dataset, but the two datasets have similar effects on foundation model pre-training. (b) Demonstration of our comprehensive benchmark for DEL. The benchmark includes a dataset of millions of data samples from 31 medical centers (DataDEL), a baseline method for comparison (MedDEL), and a new evaluation metric (NormDEL).
1. Introduction
@@
The effectiveness of foundation models depends on the abundance of pre-training data, a notion seemingly supported by consensus: more pre-training data leads to enhanced model performance. However, is this assumption truly accurate? To explore this issue, we introduce the concept of data-effective learning, which plays a significant role in the field of medical data, accelerating foundation model training and reducing storage burden in the era of big data (Figure 1(a)). The global endoscopy surgery market is undergoing significant development, with an estimated daily addition of 22,546,800,000 video frames  (end, 2023). However, a substantial presence of disruptive and invalid data (Nyeem et al., 2013) significantly hampers training efficiency and occupies a considerable amount of storage space (Yaffe, 2019). Therefore, achieving data-effective in endoscopy datasets holds the following special advantages:
@@
  • Storage Savings: Assuming the use of traditional high-definition endoscopes (1080p) (He et al., 2021b), the daily uncompressed endoscopy examination videos would require 12,756,493 TB (about 13.06 exabytes) of storage space. Condensing endoscopy datasets can result in substantial storage savings.
@@
  • Enhanced Model Efficiency: (Borgli et al., 2020) released the largest digestive system image dataset (Hyper-Kvasir) to date and showed that upon analyzing data sources, over 90% of video frames consist of disruptive and invalid data. Core critical data comprises only 2% of the entire dataset  (Jha et al., 2020, 2021b). Efficient utilization of core critical data can significantly improve the efficiency of model training, in order to achieve rapid convergence.
@@
  • Computational Resource Savings: With the use of a single RTX 3090 graphics card and the VGG16 model (Kim et al., 2016) (approximately 138 million parameters) for 32-bit floating-point (FP32) calculations, processing 325.6 images per second is achievable (Kwasny et al., 2020). In this configuration, training on the daily added video frames would require 19,200 hours. Utilizing only core critical data could save nearly 18,816 hours without compromising precision.
@@
The issue of data-effective learning is not exclusive to endoscopy datasets; it is also evident in other types of medical datasets. With the rapid expansion of future medical data, efficiently handling medical datasets is the next crucial research problem in data-driven learning methods(Rahman and Davis, 2013).
@@
In recent years, there has been relevant research on data-effective learning in natural image datasets such as adversarial samples (Feinman et al., 2017), and data biases (Wang et al., 2019), aiming to enhance the robustness and generalization capabilities of deep learning models. These works in natural datasets have shown preliminary effectiveness, indicating significant potential advantages such as data storage, computational resource savings, and efficient model training. However, despite the exponential growth of medical datasets, there is currently a lack of such exploration in the medical field (Xiao et al., 2023). This is attributed to the absence of relevant benchmarks, encompassing unified datasets, baseline methods, and comprehensive evaluation metrics, providing the academic community with a basis for developing and comparing advanced methods.
@@
In addition to the absence of benchmark, achieving data-effective in the field of endoscopy remains a huge challenge (Sharma et al., 2021) and is worth exploring in depth. Traditional data-effective techniques may struggle to work effectively with datasets that have high-dimensional features or sparse data. In endoscopy datasets, the challenge lies more in identifying subtle differences in images, which can be crucial for medical diagnosis. Therefore, accurately determining the similarity of images in endoscopy datasets is a top priority in research. Maintaining data quality and integrity is a key issue in data-effective learning (Eckerson, 2002), as it directly affects the usability and accuracy of the dataset.
Figure 2. Pipeline of the baseline method (MedDEL) for data-effective learning in our benchmark. It illustrates effective removal of disruptive and invalid data from the dataset, aiming to save storage space and computational resources while enhancing model efficiency.
@@
2. Related Work
In the natural image domain, benchmarks for assessing data-effective learning typically involve various techniques such as structural similarity index (Dosselmann and Yang, 2011), convolutional neural networks (Li et al., 2022), and local feature descriptors (Leng et al., 2018). These methods (Meyer and Bolosky, 2012; Xia et al., 2016) are evaluated on widely used image datasets like ImageNet (Deng et al., 2009; Russakovsky et al., 2015) and CIFAR (Krizhevsky et al., 2009), or specialized datasets. PXDedup (Xie et al., 2021) explores the issue of data-effective learning in JPEG images, pointing out that traditional binary stream-based techniques do not work well for compressed JPEG images.  (Chen et al., 2013) introduces a high-precision image data effective method that identifies and eliminates duplicate images through feature extraction and high-dimensional indexing (Jimenez-Rodriguez et al., 2007).  (Chen et al., 2012) uses wavelet decomposition (Mallat, 1989) to extract feature vectors from images and calculate the Manhattan distance to determine image similarity, thus achieving the detection and removal of duplicate images. However, when handling datasets with high-dimensional features (Ray et al., 2021) or sparse data (Xie et al., 2021), traditional data-effective techniques might not work effectively. MFDedup (Zou et al., 2022) surpasses existing technologies in improving data reduction rates and recovery throughput, while also reducing the cost of garbage collection. Storage and memory capacity (Cowan, 2001) can also become limiting factors.  (Harnik et al., 2012) discusses how to accurately estimate the data reduction ratio achieved through data effective and compression techniques for specific datasets. Hash methods (Periasamy and Latha, 2021) indeed have a wide application in data effective processing. The CE-Dedup (Li et al., 2021) framework combines hash-based image data-effective techniques with deep learning image classification tasks. By adjusting the deduplication threshold, effectively balances the trade-off between data reduction rate and model accuracy (Furtado and Madeira, 1999).  (Liu et al., 2016) proposes a deep supervised hash algorithm for image retrieval by providing the model with pairs of images labeled as similar/dissimilar.  (Dakshit et al., 2022) proposed a Core-set selection method based on metric explanations (CSUME) for the classification of multi-class electrocardiograms. This work offers us an effective method for selecting representative datasets within an unsupervised learning framework, which holds significant importance in the fields of medical image computing and health informatics.
@@ 2. Related Work
3. Data-Effective Learning Medical Benchmark
In this section, we introduce the definition of benchmark tasks. Data-effective learning refers to the practice of using a limited amount of data for the pre-training phase (Erhan et al., 2010). Therefore, this benchmark encourages researchers to develop advanced data-effective learning methods to generate a compact version of the originally collected large-scale dataset, thereby obtaining a compact small-scale new dataset for pre-training foundational models.
In the era of big data, large models often require massive pre-training data (Chen and Lin, 2014). However, our perspective challenges this conventional thinking, and we have undertaken the following work to explore this issue in depth.
@@ 3. Data-Effective Learning Medical Benchmark
3.1. The benchmark dataset: DataDEL
Given the urgent demand for a comprehensive benchmark in the medical field, we are facing the pressing task of integrating diverse large datasets. Currently, prevailing challenges within existing medical datasets encompass issues like the uniformity of data sources, the standardization of task execution, the limited scope of covered disease types, imbalances in dataset categories, and the uniformity of modalities (Li et al., 2023).
@@ 3.1. The benchmark dataset: DataDEL
3.2. The benchmark baseline: MedDEL
We introduce a benchmark baseline method (MedDEL) in the endoscopic medical field, which is based on the principles of the SemDeDup method (Abbas et al., 2023). Endoscopic data holds significant importance in medical diagnosis and treatment. However, due to various factors such as organ morphological changes (Smith and Nanci, 2003), lighting conditions (Clancy et al., 2012), and noise interference (Gulenko et al., 2022), the analysis of this data becomes complex and challenging.
@@ 3.2. The benchmark baseline: MedDEL
3:     for $cluster_{i}\in clusters$ do
4:         for $j=1$ to $\text{size}(cluster_{i})$ do
5:              if $\text{dis}(F_{x,j},centers_{i})>\epsilon$ then
@@
8:                  for $k=j$ to $\text{Size}(cluster_{i})$ do
9:                       if $\text{cos}(F_{j},F_{k})>\eta$ then
10:                           if $\text{dis}(F_{j},C_{i})>\text{dis}(F_{k},C_{i})$ then
11:                                Delete $I_{k}$
12:                           else
13:                                Delete $I_{j}$
14:                           end if
15:                       end if
16:                  end for
17:              end if
18:         end for
19:     end for
20:     Increment iteration counter: $t\leftarrow t+1$
21:end while
Specifically, the MedDEL method takes an image $I$ as an $i^{th}$ example, the encoder of Vision Transformer (ViT) (Dosovitskiy et al., 2020) extracts deep features from the image, ultimately producing a 768-dimensional feature output. This output serves as our information embedding. The high-dimensional representation not only captures the semantic information of the image more effectively but also addresses semantic repetitiveness issues that are challenging to resolve in low-dimensional spaces. Considering the encoding process of image $I$ through the ViT model, we can represent it with the following formula:
Table 2. Experimental setting for the parameter of $\eta$ in Algorithm  1 that controls the data remaining ratio. We set several different ratios to fully assess the performance of MedDEL under various proportions of remaining data. Additionally, we also calculate the number of training epochs required for different ratios of data at the same computational power for fair comparison.
@@
(1) $F_{x}=ViT(I)$
Where $x$ represents the final layer of the $ViT$ encoder.
However, by introducing the K-means algorithm (Macqueen, 1967), the overall computational complexity is reduced from $O(n^{2})$ to $O(N^{2}/k)$, where $k$ is the number of clusters. After applying the K-means algorithm, we obtain $k$ clusters. Assuming that the feature $F(x,i)$ belongs to the $k$ cluster, we can derive the following $k$ sequences:
(2) $C_{k}$ =$=\{F(x,i)\mid i\in N^{*},F(x,i)\in cluster_{k}\}$
@@
Figure 3. Demonstration of the feasibility of data-effective learning. We compared the performance differences between using only 5% of the pre-training data (in red) and using 100% of the data (in black) in 8 datasets. The results indicate that using only 5% of the pre-training data can achieve results comparable to using 100% of the pre-training data, which fully demonstrates the validity of the MedDEL method.
Table 3. Demonstration of the rationality of NormDEL. The table shows the performance of different proportions of pre-trained data on mIoU ($\uparrow$) as well as NormDEL ($\uparrow$) in eight different datasets. The results indicate that the performance of mIoU is almost comparable in different scales, suggesting that it primarily measures performance without considering the scale of the pre-training data used. In contrast, NormDEL incorporates the factor of data scale, and can evaluate data compactness even with only 5% of the data used. This demonstrates the rationality of the NormDEL method, as it not only assesses performance but also effectively utilizes the data scale.
@@
3.3. The benchmark evaluation metric: NormDEL
Currently, there is no comprehensive objective metric for evaluating the performance of data-effective learning algorithms, where comprehensive refers to integrating both the test accuracy of downstream tasks and the compactness of pre-training data. Hence, We propose a new data-effective metric, the Normalized Data-Effective Learning Index (NormDEL). Our objective is that within the same task, if a model can achieve the same level of performance in downstream tasks using fewer pre-training data, its data-effective performance should be considered superior. Taking the commonly used segmentation task in endoscopy as an example, where the general metric for measuring performance is often mIoU (mean Intersection over Union), we can formulate the following equation:
(4) $DEL=mIoU\cdot e^{-\alpha\cdot R}$